### v1

# Private Heterogeneous Federated Learning Without a Trusted Server Revisited: Error-Optimal and Communication-Efficient Algorithms for Convex Losses

Abstract
We revisit the problem of federated learning (FL) with private data from people who do not trust the server or other silos/clients. In this context, every silo (e.g. hospital) has data from several people (e.g. patients) and needs to protect the privacy of each person’s data (e.g. health records), even if the server and/or other silos try to uncover this data. Inter-Silo Record-Level Differential Privacy (ISRL-DP) prevents each silo’s data from being leaked, by requiring that silo $i$’s communications satisfy item-level differential privacy. Prior work Lowy and Razaviyayn (2023a) characterized the optimal excess risk bounds for ISRL-DP algorithms with homogeneous (i.i.d.) silo data and convex loss functions. However, two important questions were left open: (1) Can the same excess risk bounds be achieved with heterogeneous (non-i.i.d.) silo data? (2) Can the optimal risk bounds be achieved with fewer communication rounds? In this paper, we give positive answers to both questions. We provide novel ISRL-DP FL algorithms that achieve the optimal excess risk bounds in the presence of heterogeneous silo data. Moreover, our algorithms are more communication-efficient than the prior state-of-the-art. For smooth loss functions, our algorithm achieves the optimal excess risk bound and has communication complexity that matches the non-private lower bound. Additionally, our algorithms are more computationally efficient than the previous state-of-the-art.
1 Introduction
Federated learning (FL) is a distributed machine learning paradigm in which multiple silos (a.k.a. clients), such as hospitals or cell-phone users, collaborate to train a global model. In FL, silos store their data locally and exchange focused updates (e.g. stochastic gradients), sometimes making use of a central server (Kairouz et al., 2021). FL has been applied in myriad domains, from consumer digital products (e.g. Google’s mobile keyboard Hard et al. (2018) and Apple’s iOS Apple (2019)) to healthcare Courtiol et al. (2019), finance FedAI (2019), and large language models (LLMs) Hilmkil et al. (2021).
One of the primary reasons for the introduction of FL was to enhance protection of the privacy of people’s data (McMahan et al., 2017). Unfortunately, local storage is not sufficient to prevent data from being leaked, because the model parameters and updates communicated between the silos and the central server can reveal sensitive information (Zhu and Han, 2020; Gupta et al., 2022). For example, Gupta et al. (2022) attacked a FL model for training LLMs to uncover private text data.
Differential privacy (DP) Dwork et al. (2006) guarantees that private data cannot be leaked. Different variations of DP have been considered for FL. Central DP prevents the final trained FL model from leaking data to an external adversary Jayaraman and Wang (2018); Noble et al. (2022). However, central DP has two major drawbacks: (1) it does not provide a privacy guarantee for each individual silo; and (2) it does not ensure privacy when an attacker/eavesdropper has access to the server or to another silo.
Another notion of DP for FL is user-level DP McMahan et al. (2018); Geyer et al. (2017); Levy et al. (2021). User-level DP mitigates the drawback (1) of central DP, by providing privacy for the complete local data set of each silo. User-level DP is practical in the cross-device FL setting, where each silo is a single person (e.g. cell-phone user) with a large number of records (e.g. text messages). However, user-level DP still permits privacy breaches if an adversary has access to the server or eavesdrops on the communications between silos. Moreover, user-level DP is not well-suited for cross-silo federated learning, where silos represent organizations like hospitals, banks, or schools that house data from many individuals (e.g. patients, customers, or students). In the cross-silo FL context, each person possesses a record, referred to as an “item,” which may include sensitive data. Therefore, an appropriate notion of DP for cross-silo FL should safeguard the privacy of each individual record (i.e. “item-level differential privacy”) within silo $i$, rather than the complete aggregated data of silo $i$.
Following Lowy and Razaviyayn (2023a); Lowy et al. (2023a); Liu et al. (2022) (among others), this work considers inter-silo record-level DP (ISRL-DP). ISRL-DP requires that the full transcript of messages sent by silo $i$ satisfy item-level DP, for all silos $i$. Thus, ISRL-DP guarantees the privacy of each silo’s local data, even in the presence of an adversary with access to the server, other silos, or the communication channels between silos and the server. See Figure 1 for an illustration. If each silo only has one record, then ISRL-DP reduces to local DP Kasiviswanathan et al. (2011); Duchi et al. (2013). However, in the FL setting where each silo has many records, ISRL-DP FL is more appropriate and permits much higher accuracy than local DP Lowy and Razaviyayn (2023a). Moreover, ISRL-DP implies user-level DP if the ISRL-DP parameters are sufficiently small Lowy and Razaviyayn (2023a). Thus, ISRL-DP is a practical privacy notion for cross-silo and cross-device FL when individuals do not trust the server or other silos/clients/devices with their sensitive data.
Figure 1: ISRL-DP maintains the privacy of each patient’s record, provided the patient’s own hospital is trusted. Silo $i$’s messages are item-level DP, preventing data leakage, even if the server/other silos collude to decode the data of silo $i$.
Problem Setup.
Consider a FL environment with $N$ silos, where each silo has a local data set with $n$ samples: $X_{i}=(x_{i,1},\cdots,x_{i,n})$ for $i\in[N]:=\{1,\ldots,N\}$. At the beginning of every communication round $r$, silos receive the global model $w_{r}$. Silos then utilize their local data to enhance the model, and send local updates to the server (or to other silos for peer-to-peer FL). The server (or other silos) uses the local silo updates to update the global model.
For each silo $i$, let $\mathcal{D}_{i}$ be an unknown probability distribution on a data domain $\mathcal{X}$. Let $f:\mathcal{W}\times\mathcal{X}\to\mathbb{R}$ be a loss function (e.g. cross-entropy loss for logistic regression), where $\mathcal{W}\subset\mathbb{R}^{d}$ is a parameter domain. Assume $f(\cdot,x)$ is convex. Silo $i$’s local objective is to minimize its expected population/test loss, which is
$F_{i}(w):=\mathbb{E}_{x_{i}\sim\mathcal{D}_{i}}[f(w,x_{i})].$ (1)
Our (global) objective is to find a model that achieves small error for all silos, by solving the FL problem
$\min_{w\in\mathcal{W}}\left\{F(w):=\frac{1}{N}\sum_{i=1}^{N}F_{i}(w)\right\},$ (FL)
while maintaining the privacy of each silo’s local data under ISRL-DP.
Assume that the samples $\{x_{i,j}\}_{i\in[N],j\in[n]}$ are independent. The FL problem is homogeneous if the data is i.i.d., i.e., $\mathcal{D}_{i}=\mathcal{D}_{j}$ for all $i,j\in[N]$. In this work, we focus on the more challenging heterogeneous (non-i.i.d.) case: i.e. the data distributions $\{\mathcal{D}_{i}\}_{i=1}^{N}$ may be arbitrary. This case arises in many practical FL settings Kairouz et al. (2021).
Contributions.
The quality of a private algorithm $\mathcal{A}$ for solving Problem FL is measured by its excess risk (a.k.a. excess population/test loss), $\mathbb{E}[F(\mathcal{A}(\mathbf{X}))]-F^{*}$, where $F^{*}=\inf_{w\in\mathcal{W}}F(w)$ and the expectation is over the random draw of the data $\mathbf{X}=(X_{1},\ldots,X_{N})$ as well as the randomness of $\mathcal{A}$Lowy and Razaviyayn (2023a) characterized the minimax optimal excess risk (up to logarithms) for convex ISRL-DP FL in the homogeneous (i.i.d.) case:
$\widetilde{\Theta}\left(\frac{1}{\sqrt{N}}\left(\frac{1}{\sqrt{n}}+\frac{\sqrt{d\log(1/\delta)}}{\varepsilon n}\right)\right),$ (2)
where $d$ is the dimension of the parameter space and $\varepsilon,\delta$ are the privacy parameters. For heterogeneous FL, the state-of-the-art excess risk bound is $O\left(1/\sqrt{Nn}+\left(\sqrt{d\log(1/\delta)}/(\varepsilon n\sqrt{N})\right)^{4/5}\right)$, assuming smoothness of the loss function Lowy and Razaviyayn (2023a). Lowy and Razaviyayn (2023a) left the following question open:
Question 1. Can the optimal ISRL-DP excess risk in (2) for solving Problem FL be attained in the presence of heterogeneous silo data?
Contribution 1. We give a positive answer to Question 1. Moreover, we do not require smoothness of the loss function to attain the optimal rate: see Theorem 3.1.
In practical FL settings (especially cross-device settings), it can be expensive or slow for silos to communicate with the server or with each other Kairouz et al. (2021). Thus, another important desiderata for FL algorithms is communication efficiency, which is measured by communication complexity: the number of communication rounds $R$ that are required to achieve excess risk $\alpha$. In the homogeneous setting, the state-of-the-art communication complexity for an algorithm achieving optimal excess risk is due to Lowy and Razaviyayn (2023a, Theorem D.1):
$R_{\text{SOTA}}=\widetilde{O}\left(\min\left\{Nn,\frac{Nn^{2}\varepsilon^{2}}{d}\right\}\right).$ (3)
Question 2. Can the optimal ISRL-DP excess risk in (2) for solving Problem FL be attained in fewer communication rounds $R$, for $R\ll R_{SOTA}$?
Contribution 2. We answer Question 2 positively. For smooth losses, our algorithm achieves optimal excess risk with significantly improved communication complexity:
$R_{\text{smooth}}=\widetilde{O}\left(\min\left\{(Nn)^{1/4},\frac{N^{1/4}n^{1/2}\varepsilon^{1/2}}{d^{1/4}}\right\}\right).$ (4)
Our communication complexity in (4) matches the non-private lower bound of Woodworth et al. (2020) in the high heterogeneity regime (Theorem 2.4), hinting at the communication-optimality of our algorithm.
For non-smooth loss functions, our communication complexity is
$R_{\text{non-smooth}}=\widetilde{O}\left(\min\left\{(Nn)^{1/2},\frac{N^{1/2}n\varepsilon}{d^{1/2}}\right\}\right),$ (5)
a major improvement over the prior state-of-the-art bound in (3).
Moreover, we achieve these improved communication complexity and optimal excess risk bounds without assuming homogeneity of silo data.
When $N=1$ in (FL), ISRL-DP FL reduces to central DP stochastic convex optimization (SCO), which has been studied extensively Bassily et al. (2019); Feldman et al. (2020); Asi et al. (2021); Zhang et al. (2022). In this centralized setting, communication complexity is usually referred to as iteration complexity. Even in the special case of $N=1$, our iteration complexity for smooth losses improves over the prior state-of-the-art result Zhang et al. (2022).
Another important property of FL algorithms is computational efficiency, which we measure by (sub)gradient complexity: the number of stochastic (sub)gradients $T$ that an algorithm must compute to achieve excess risk $\alpha$.
In the homogeneous case, the state-of-the-art gradient complexity for an ISRL-DP FL algorithm that attains optimal excess risk is due to Lowy and Razaviyayn (2023a, Theorem D.1). For smooth losses and $\varepsilon=\Theta(1)$, Lowy et al. (2023a) obtained the follow gradient complexity
$T_{\text{SOTA}}\!=\!\widetilde{O}\left(N^{2}\min\left\{n,\frac{n^{2}}{d}\right\}+N^{3/2}\min\left\{n^{3/2},\frac{n^{2}}{d^{1/2}}\right\}\right).$ (6)
Question 3. Can the optimal ISRL-DP excess risk in (2) for solving Problem FL be attained with smaller gradient complexity $T\ll T_{\text{SOTA}}$?
Contribution 3. We give a positive answer to Question 3. We provide an ISRL-DP FL algorithm with optimal excess risk and improved gradient complexity (see Theorem 2.1). When $d=\Theta(n)$ and $\varepsilon=\Theta(1)$, our gradient complexity bound simplifies to
$T_{\text{smooth}}=\widetilde{O}\left(N^{5/4}n^{1/4}+(Nn)^{9/8}\right).$ (7)
In Theorem 3.4, we also improve over the state-of-the-art subgradient complexity bound of Lowy and Razaviyayn (2023a, Theorem D.2) for non-smooth losses. Moreover, in contrast to these earlier results, we do not assume homogeneous data.
We summarize our main results in Table 1 and Table 2.
Table 1: Comparison vs. SOTA for Smooth Loss Functions. In Tables 1-2, [LR’23] refers to Lowy and Razaviyayn (2023a); we omit logs and fix $M=N$, $d=\Theta(n)$, $\varepsilon=\Theta(1)$, $L=\beta=D=1$. See theorems and the appendix for more general cases.
Algorithm & Setting Excess Risk Communication Complexity Gradient Complexity
[LR’23] Alg. 2 (i.i.d.) optimal $Nn$ $N^{2}n^{2}$
[LR’23] Alg. 1 (non-i.i.d.) suboptimal $N^{1/5}n^{1/5}$ $Nn$
Alg. 1 (non-i.i.d.) optimal $N^{1/4}n^{1/4}$ $N^{5/4}n^{1/4}+(Nn)^{9/8}$
Table 2: Comparison vs. SOTA for Nonsmooth Loss Functions. First 3 algorithms use Nesterov smoothing.
Algorithm & Setting Excess Risk Communication Complexity Gradient Complexity
[LR’23] Alg. 2 (i.i.d.) optimal $Nn$ $N^{2}n^{2}$
[LR’23] Alg. 1 (non-i.i.d.) suboptimal $N^{1/3}n^{1/3}$ $Nn$
Alg. 1 (non-i.i.d.) optimal $N^{1/2}n^{1/2}$ Omitted (see Remark 3.3)
Alg. 4 (non-i.i.d.) optimal $Nn$ $N^{2}n+N^{3/2}n^{3/2}$
Algorithm 6 (non-i.i.d.) optimal $Nn^{5/4}$ $N^{3/2}n^{3/4}+N^{5/4}n^{11/8}$
Our Algorithms and Techniques.
Our algorithms combine and extend various private optimization techniques in novel ways.
Our ISRL-DP Algorithm 1 for smooth FL builds on the ISRL-DP Accelerated MB-SGD approach used by Lowy and Razaviyayn (2023a) for federated empirical risk minimization (ERM). We call this algorithm repeatedly to iteratively solve a carefully chosen sequence of regularized ERM subproblems, using the localization technique of Feldman et al. (2020). We obtain our novel optimal excess risk bounds for heterogeneous FL with an algorithmic stability Bousquet and Elisseeff (2002) argument. We make the key observation that the stability and generalization of regularized ERM in Shalev-Shwartz et al. (2009) does not require the data to be identically distributed. By combining this observation with a bound for ISRL-DP Accelerated MB-SGD, we obtain our excess risk bound.
It is not immediately clear that the approach just described should work, because the iterates of Accelerated MB-SGD are not necessarily stable. The instability of acceleration may explain why Lowy and Razaviyayn (2023a) used ISRL-DP Accelerated MB-SGD only for ERM and not for minimizing the population risk. We overcome this obstacle with an alternative analysis that leverages the stability of regularized ERM, the convergence of ISRL-DP Accelerated MB-SGD to an approximate minimizer of the regularized empirical loss, and localization. This argument enables us to show that our algorithm has optimal excess risk. By carefully choosing algorithmic parameters, we also obtain state-of-the-art communication and gradient complexities.
To extend our algorithm to the non-smooth case, we use two different techniques. One is Nesterov smoothing Nesterov (2005), which results in an algorithm with favorable communication complexity. The other approach is to replace the ISRL-DP Accelerated MB-SGD ERM subsolver by an ISRL-DP Subgradient Method. This technique yields an algorithm with favorable (sub)gradient complexity. Third, we use randomized convolution smoothing, as in Kulkarni et al. (2021), to obtain another favorable gradient complexity bound.
1.1 Preliminaries
Differential Privacy.
Let $\mathbb{X}=\mathcal{X}^{n\times N}$ and $\rho:\mathbb{X}^{2}\to[0,\infty)$ be a distance between distributed data sets. Two distributed data sets $\mathbf{X},\mathbf{X}^{\prime}\in\mathbb{X}$ are $\rho$-adjacent if $\rho(\mathbf{X},\mathbf{X}^{\prime})\leq 1$. Differential privacy (DP) prevents an adversary from distinguishing between the outputs of algorithm $\mathcal{A}$ when it is run on adjacent databases:
Definition 1.1 (Differential Privacy Dwork et al. (2006)).
Let $\varepsilon\geq 0,~{}\delta\in[0,1)$. A randomized algorithm $\mathcal{A}:\mathbb{X}\to\mathcal{W}$ is $(\varepsilon,\delta)$-differentially private (DP) for all $\rho$-adjacent data sets $\mathbf{X},\mathbf{X}^{\prime}\in\mathbb{X}$ and all measurable subsets $S\subseteq\mathcal{W}$, we have
$\mathbb{P}(\mathcal{A}(\mathbf{X})\in S)\leq e^{\varepsilon}\mathbb{P}(\mathcal{A}(\mathbf{X}^{\prime})\in S)+\delta.$ (8)
Definition 1.2 (Inter-Silo Record-Level Differential Privacy).
Let $\rho:\mathcal{X}^{2n}\to[0,\infty)$, $\rho(X_{i},X^{\prime}_{i}):=\sum_{j=1}^{n}\mathbbm{1}_{\{x_{i,j}\neq x_{i,j}^{\prime}\}}$, $i\in[N]$. A randomized algorithm $\mathcal{A}$ is $(\varepsilon,\delta)$-ISRL-DP if for all $i\in[N]$ and all $\rho$-adjacent silo data sets $X_{i},X^{\prime}_{i}$, the full transcript of silo $i$’s sent messages satisfies (8) for any fixed settings of other silos’ data.
Notation and Assumptions.
[…]
We write $a\lesssim b$ if $\exists\,C>0$ such that $a\leq Cb$. We use $a=\widetilde{O}(b)$ to hide poly-logarithmic factors.
We assume the following throughout:
Assumption 1.3.
  1. 1.
    $\mathcal{W}\subset\mathbb{R}^{d}$ is closed, convex. We assume $\|w-w^{\prime}\|\leq D$, $\forall w,w^{\prime}\in\mathcal{W}$.
  2. 2.
    $f(\cdot,x)$ is $L$-Lipschitz and convex for all $x\in\mathcal{X}$. In some places, we assume that $f(\cdot,x)$ is $\beta$-smooth.
  3. 3.
    In each round $r$, a uniformly random subset $S_{r}\subset[N]$ of $M$ silos is available to communicate with the server.
For simplicity, in the main body, we often assume $M=N$. The Appendix contains the general statements of all results, complete proofs, and a further discussion of related work.
2 Localized ISRL-DP Accelerated MB-SGD for Smooth Losses
We start with the smooth case. Combining iterative localization techniques of Feldman et al. (2020); Asi et al. (2021) with the multi-stage ISRL-DP Accelerated MB-SGD of Lowy and Razaviyayn (2023a), our proposed Algorithm 1 achieves optimal excess risk and state-of-the-art communication complexity for heterogeneous FL under ISRL-DP.
Algorithm 1 Localized ISRL-DP Accelerated MB-SGD
1:Datasets $X_{l}\in\mathcal{X}^{n}$ for $l\in[N]$, loss function $f$, constraint set $\mathcal{W}$, initial point $w_{0}$, subroutine parameters $\{R_{i}\}_{i=1}^{\lfloor\log_{2}n\rfloor}\subset\mathbb{N}$, $\{K_{i}\}_{i=1}^{\lfloor\log_{2}n\rfloor}\subset[n]$.
2:Set $\tau=\lfloor\log_{2}n\rfloor$.
3:Set $p=\max(\tfrac{1}{2}\log_{n}(M)+1,3)$
4:for $i=1$ to $\tau$ do
5:     Set $\lambda_{i}=\lambda\cdot 2^{(i-1)p}$, $n_{i}=\lfloor n/2^{i}\rfloor$, $D_{i}=2L/\lambda_{i}$.
6:     Each silo $l\in[N]$ draws disjoint batch $B_{i,l}$ of $n_{i}$ samples from $X_{l}$.
7:     Let $\hat{F}_{i}(w)=\frac{1}{n_{i}N}\sum_{l=1}^{N}\sum_{x_{l,j}\in B_{i,l}}f(w;x_{l,j})+\frac{\lambda_{i}}{2}\lVert w-w_{i-1}\rVert^{2}$.
8:     Call the multi-stage $(\varepsilon,\delta)$-ISRL-DP implementation of Algorithm 2 with loss function $\hat{F}_{i}(w)$, data $X_{l}=B_{i,l}$, $R=R_{i}$, $K=K_{i}$, initialization $w_{i-1}$, constraint set $\mathcal{W}_{i}=\{w\in\mathcal{W}:\lVert w-w_{i-1}\rVert\leq D_{i}\}$, and $\mu=\lambda_{i}$. Denote the output by $w_{i}$.
9:end for
10:return the last iterate $w_{\tau}$
Algorithm 2 Accelerated ISRL-DP MB-SGD Lowy and Razaviyayn (2023a)
1:Datasets $X_{l}\in\mathcal{X}^{n}$ for $l\in[N]$, loss function $\hat{F}(w)=\frac{1}{nN}\sum_{l=1}^{N}\sum_{x\in X_{l}}f(w,x)$, constraint set $\mathcal{W}$, initial point $w_{0}$, strong convexity modulus $\mu\geq 0$, privacy parameteres $(\varepsilon,\delta)$, iteration count $R\in\mathbb{N}$, batch size $K\in[n]$, step size parameters $\{\eta_{r}\}_{r\in[R]},\{\alpha_{r}\}_{r\in[R]}$ specified in Section 7.
2:Initialize $w_{0}^{ag}=w_{0}\in\mathcal{W}$ and $r=1$.
3:for $r\in[R]$ do
4:     Server updates and broadcasts
5:$w_{r}^{md}=\frac{(1-\alpha_{r})(\mu+\eta_{r})}{\eta_{r}+(1-\alpha_{r}^{2})\mu}w_{r-1}^{ag}+\frac{\alpha_{r}[(1-\alpha_{r})\mu+\eta_{r}]}{\eta_{r}+(1-\alpha_{r}^{2})\mu}w_{r-1}$
6:     for $i\in S_{r}$ in parallel do
7:         Silo $i$ draws $\{x_{i,j}^{r}\}_{j=1}^{K}$ from $X_{i}$ (with replacement) and privacy noise $u_{i}\sim\mathcal{N}(0,\sigma^{2}\mathbf{I}_{d})$ for proper $\sigma^{2}$.
8:         Silo $i$ computes $\widetilde{g}_{r}^{i}:=\frac{1}{K}\sum_{j=1}^{K}\nabla f(w_{r}^{md},x_{i,j}^{r})+u_{i}$.
9:     end for
10:     Server aggregates $\widetilde{g}_{r}:=\frac{1}{M}\sum_{i\in S_{r}}\widetilde{g}_{r}^{i}$ and updates:
11:     $w_{r}:=\operatorname*{arg\,min}_{w\in\mathcal{W}}\left\{\alpha_{r}\left[\langle\widetilde{g}_{r},w\rangle+\frac{\mu}{2}\|w_{r}^{md}-w\|^{2}\right]\right.$ $\left.+\left[(1-\alpha_{r})\frac{\mu}{2}+\frac{\eta_{r}}{2}\right]\|w_{r-1}-w\|^{2}\right\}$.
12:     Server updates and broadcasts
13:$w_{r}^{ag}=\alpha_{r}w_{r}+(1-\alpha_{r})w_{r-1}^{ag}.$
14:end for
15:
16:return: $w_{R}^{ag}$.
We recall ISRL-DP Accelerated MB-SGD in Algorithm 2. It is a distributed, ISRL-DP version of the AC-SA algorithm Ghadimi and Lan (2012). For strongly convex losses, the multi-stage implementation of ISRL-DP Accelerated MB-SGD, given in Algorithm 5 in Section 7, offers improved excess risk Lowy and Razaviyayn (2023a). Algorithm 5 is a distributed, ISRL-DP version of the multi-stage AC-SA Ghadimi and Lan (2013).
Building on Algorithm 5, we describe our algorithm as follows (see Algorithm 1 for pseudocode). The distributed learning process is divided into $\tau=\lfloor\log_{2}n\rfloor$ phases. In each phase $i\in[\tau]$, all silos work together (via communication with the central server) to iteratively solve a regularized ERM problem with ISRL-DP. The regularized ERM problem in phase $i$ is defined over $N$ local batches of data, each containing $n_{i}$ disjoint samples (cf. $\hat{F}_{i}(w)$ in Line 7). To find the approximate constrained minimizer of $\hat{F}_{i}(w)$ privately, we apply Algorithm 5 for a careful choice of the number of rounds $R_{i}$ and the batch size $K_{i}\in[n_{i}]$ (cf. Line 8). The output $w_{i}$ of phase $i$ affects phase $i+1$ in three ways: (i) regularization center used to define $\hat{F}_{i+1}$; (ii) initialization for the next call of Algorithm 5; and (iii) constraint set $\mathcal{W}_{i+1}$. We enforce stability (hence generalization) of our algorithm via regularization and localization: e.g., as $i$ increases, we increase the regularization parameter $\lambda_{i}$ to prevent $w_{i+1}$ from moving too far away from $w_{i}$ and $w^{*}$.
The following theorem captures our main results of this section: Algorithm 1 can achieve the optimal excess risk, regardless of the heterogeneity, in a communication-efficient and gradient-efficient manner.
Theorem 2.1 (Upper Bound for Smooth Losses).
Let $f(\cdot,x)$ be $\beta$-smooth and $M=N$. Assume $\varepsilon\leq 2\ln(2/\delta),\delta\in(0,1)$. Then, there exist parameter choices such that Algorithm 1 is $(\varepsilon,\delta)$-ISRL-DP and has the following excess risk
$\!\!\mathbb{E}F(w_{\tau})\!-\!F(w^{*})\!=\!\widetilde{O}\!\left(\frac{LD}{\sqrt{N}}\!\left(\frac{1}{\sqrt{n}}\!+\!\frac{\sqrt{d\log(1/\delta)}}{\varepsilon n}\right)\!\!\right).$ (9)
Moreover, the communication complexity of Algorithm 1 is
$\widetilde{O}\left(\frac{\sqrt{\beta D}N^{1/4}}{\sqrt{L}}\left(\min\left\{\sqrt{n},\frac{\varepsilon n}{\sqrt{d\ln(1/\delta)}}\right\}\right)^{1/2}+1\right).$
Assuming $d=\Theta(n)$ and $\varepsilon=\Theta(1)$, the gradient complexity of Algorithm 1 is
$\widetilde{O}\left(N^{5/4}n^{1/4}(\beta D/L)^{1/2}+Nn+(Nn)^{9/8}(\beta D/L)^{1/4}\right).$
For general $d,n,\varepsilon$, the gradient complexity expression is complicated, and is given in the Appendix 8.1.
Remark 2.2 (Optimal risk in non-i.i.d private FL).
The excess risk bound in (9) matches the optimal i.i.d risk bound up to log factors (cf. Theorem 2.2 in Lowy and Razaviyayn (2023a)), even when silo data is arbitrarily heterogeneous across silos. To the best of our knowledge, our algorithm is the first to have this property, resolving an open question of Lowy and Razaviyayn (2023a). The prior state-of-the-art bound in Lowy and Razaviyayn (2023a, Theorem 3.1) is suboptimal by a factor of $\widetilde{O}((\sqrt{d}/(\varepsilon n\sqrt{N}))^{1/5})$.
Remark 2.3 (Improved communication and gradient complexity).
The communication and gradient complexities of our Algorithm 1 significantly improve over the previous state-of-the-art for ISRL-DP FL: recall (3), (4), (6), and (7).
Our algorithm has state-of-the-art communication complexity, even in the simple case of $N=1$, where ISRL-DP FL reduces to central DP SCO. In fact, the prior state-of-the-art iteration complexity bound for DP SCO was $O((\beta D/L)\min\{\sqrt{n},\varepsilon n/\sqrt{d\ln(1/\delta)}\})$ Zhang et al. (2022). By comparison, when $N=1$, our algorithm’s communication complexity is the square root of this bound. Note that when $N=1$, our algorithm is essentially the same as the algorithm of Kulkarni et al. (2021), but we do not incorporate convolution smoothing here since we are assuming smoothness of the loss.
We can also compare our gradient complexity results against the state-of-the-art central DP SCO algorithms when $N=1$. As an illustration, consider the interesting regime $\varepsilon=\Theta(1)$ and $d=\Theta(n)$. For smooth and convex losses, when (i) $\beta\leq O(\sqrt{n}L/D)$, algorithms in both Feldman et al. (2020) and Zhang et al. (2022) achieve optimal risk using $\widetilde{O}(n)$ gradient complexity. For (ii) $\beta\geq\Omega(\sqrt{n}L/D)$, the two algorithms proposed by Feldman et al. (2020) fail to guarantee optimal risk, whereas (Zhang et al., 2022, Algorithm 2) continues to attain optimal risk with gradient complexity $\widetilde{O}(n^{3/4}\sqrt{\beta D/L})$. By comparison with these results, for case (i), our Algorithm 1 achieves optimal risk with gradient complexity $\widetilde{O}(n^{5/4})$. For case (ii), as in (Zhang et al., 2022), our Algorithm 1 continues to achieve optimal risk with gradient complexity $\widetilde{O}(n^{9/8}(\beta D/L)^{1/4}+n^{1/4}(\beta D/L)^{1/2}+n)$. Thus, our algorithm is faster than the state-of-the-art result of Zhang et al. (2022) when $\beta D/L\geq n^{3/2}$. In the complementary parameter regime, however, the algorithm of Zhang et al. (2022) (which is not ISRL-DP) is faster. We discuss possible ways to close this gap in Section 5.
Comparison with Non-Private Communication Complexity Lower Bound.
Theorem 2.1 trivially extends to the unconstrained optimization setting in which $D=\|w_{0}-w^{*}\|$ for $w^{*}\in\operatorname*{arg\,min}_{w\in\mathbb{R}^{d}}F(w)$. Moreover, our excess risk bound is still optimal for the unconstrained case: a matching lower bound is obtained by combining the technique for unconstrained DP lower bounds in Liu and Lu (2021) with the constrained ISRL-DP FL lower bounds of Lowy and Razaviyayn (2023a).
Let us compare our communication complexity upper bound against the non-private communication complexity lower bound of Woodworth et al. (2020). Define the following parameter which describes the heterogeneity of the silos at the optimum $w^{*}=\operatorname*{arg\,min}_{w\in\mathbb{R}^{d}}F(w)$:
$\zeta_{*}^{2}=\frac{1}{N}\sum_{i=1}^{N}\|\nabla F_{i}(w^{*})\|^{2}.$
The lower bound holds for the class of distributed zero-respecting algorithms (defined in Appendix 11), which includes most non-private FL algorithms, such as MB-SGD, Accelerated MB-SGD, local SGD/FedAvg, and so on.
Theorem 2.4 (Communication Lower Bound Woodworth et al. (2020)).
Fix $M=N$ and suppose $\mathcal{A}$ is a distributed zero-respecting algorithm with excess risk
$\mathbb{E}F(\mathcal{A}(X))-F^{*}\lesssim\frac{LD}{\sqrt{N}}\left(\frac{1}{\sqrt{n}}+\frac{\sqrt{d\ln(1/\delta)}}{\varepsilon n}\right)$
in $\leq R$ rounds of communications on any $\beta$-smooth FL problem with heterogeneity $\zeta_{*}$ and $\|w_{0}-w^{*}\|\leq D$. Then,
$R$ $\gtrsim N^{1/4}\left(\min\left\{\sqrt{n},\frac{\varepsilon n}{\sqrt{d\ln(1/\delta)}}\right\}\right)^{1/2}$
×$\;\;\;\;\times\min\left(\frac{\sqrt{\beta D}}{\sqrt{L}},\frac{\zeta_{*}}{\sqrt{\beta LD}}\right).$
Remark 2.5.
Our communication complexity in Theorem 2.1 matches the lower bound on the communication cost in Theorem 2.4 when $\zeta_{*}\gtrsim\beta D$ (i.e., high heterogeneity).
There are several reasons why we cannot quite assert that Theorem 2.4 implies that Algorithm 1 is communication-optimal. First, the lower bound does not hold for randomized algorithms that are not zero-respecting (e.g. our ISRL-DP algorithms). However, Woodworth et al. (2020) note that the lower bound should be extendable to all randomized algorithms by using the random rotations techniques of Woodworth and Srebro (2016); Carmon et al. (2020). Second, the lower bound construction of Woodworth et al. (2020) is not $L$-Lipschitz. However, we believe that their quadratic construction can be approximated by a Lipschitz function (e.g. by using an appropriate Huber function). Third, as is standard in non-private complexity lower bounds, the construction requires the dimension $d$ to grow with $R$. This third issue seems challenging to overcome, and may require a fundamentally different proof approach. Thus, a rigorous proof of a communication complexity lower bound for ISRL-DP algorithms and Lipschitz functions is an interesting topic for future work.
Sketch of the Proof of Theorem 2.1.
We end this section with a high-level overview of the key steps needed to establish Theorem 2.1.
Proof sketch.
Privacy: We choose parameters so that each call to Algorithm 5 is $(\varepsilon,\delta)$-ISRL-DP. Then, the full Algorithm 1 is $(\varepsilon,\delta)$-ISRL-DP by parallel composition, since silo $l$ samples disjoint local batches across phases ($\forall l\in[N]$).
Excess risk: Following Feldman et al. (2020); Asi et al. (2021), we start with the following error decomposition: Let $\hat{w}_{0}=w^{*}$ for analysis only, and write
$\mathbb{E}[F(w_{\tau})]-F(w^{*})$ $=\mathbb{E}[F(w_{\tau})-F(\hat{w}_{\tau})]$ (10)
$+\sum_{i=1}^{\tau}\mathbb{E}[F(\hat{w}_{i})-F(\hat{w}_{i-1})],$ (11)
where $\hat{w}_{i}=\operatorname*{arg\,min}_{w\in\mathcal{W}}\hat{F}_{i}(w)$ for $i\in[\tau]$. Then, it remains to bound (10) and (11), respectively. To this end, we first show that $w_{i}$ in Algorithm 1 is close to $\hat{w}_{i}$ for all $i\in[\tau]$. Here, we mainly leverage the excess empirical risk bound of Algorithm 5 for ERM with strongly convex and smooth loss functions. In particular, by the empirical risk bound in Lowy and Razaviyayn (2023a, Theorem F.1) and $\lambda_{i}$-strongly convex of $\hat{F}_{i}$, we can show that the following bound holds for all $i\in[\tau]$:
$\mathbb{E}\left[\left\|w_{i}-\hat{w}_{i}\right\|^{2}\right]\leq\widetilde{O}\left(\frac{L^{2}}{\lambda_{i}^{2}N}\cdot\frac{d\log(1/\delta)}{n_{i}^{2}\varepsilon^{2}}\right).$ (12)
We bound (10) via (12), Lipschitzness, and Jensen’s Inequality.
To bound (11), we first leverage a key observation that stability and generalization of the empirical minimizer of strongly-convex loss does not require homogeneous data. This observation that allows us to handle the non-i.i.d. case optimally. Specifically, by the stability result, we have
$\mathbb{E}[F(\hat{w}_{i})\!-\!F(\hat{w}_{i-1})]\lesssim\frac{\lambda_{i}\mathbb{E}[\|\hat{w}_{i-1}-w_{i-1}\|^{2}]}{2}\!+\!\frac{L^{2}}{\lambda_{i}n_{i}M}.$
We obtain a bounds on (11) by combining the above inequality with (12). Finally, by putting everything together and leveraging the geometrical schedule of $\lambda_{i},n_{i}$, we obtain the final excess population risk bound. ∎
An alternative proof approach would be to try to show that Algorithm 5 is stable directly, e.g. by using the tools that Hardt et al. (2016) use for proving stability of SGD. However, this approach seems unlikely to yield the same tight bounds that we obtain, since acceleration may impede stability of the iterates. Instead, we establish our stability and generalization guarantee indirectly: We use the facts that reguralized ERM is stable and that Algorithm 5 approximates regularized ERM.
3 Error-Optimal Heterogeneous ISRL-DP FL for Non-Smooth Losses
In this section, we turn to the case of non-smooth loss functions. We modify Algorithm 1 to obtain two algorithms that can handle the non-smooth case. Our algorithms are the first to achieve the optimal excess risk for heterogeneous ISRL-DP FL with non-smooth loss functions.
3.1 Error-Optimal and Communication-Efficient ISRL-DP FL for Non-Smooth Losses
Our first algorithm is based on the technique of Nesterov smoothing Nesterov (2005): For a non-smooth function $f$, Moreau-Yosida regularization is used to approximate $f$ by the $\beta$-smooth $\beta$-Moreau envelope
$f_{\beta}(w):=\min_{v\in\mathcal{W}}\left(f(v)+\frac{\beta}{2}\|w-v\|^{2}\right).$
We then optimize this smooth function using our Algorithm 1. See Appendix 10 for details.
Theorem 3.1 (Non-smooth FL via smoothing).
Let $M=N$. Then the combination of Algorithm 1 with Nesterov smoothing yields an $(\varepsilon,\delta)$-ISRL-DP algorithm with optimal excess population risk as in (9). The communication complexity is
$\widetilde{O}\left(\sqrt{N}\min\left\{\sqrt{n},\frac{\varepsilon n}{\sqrt{d\ln(1/\delta)}}\right\}+1\right).$
Remark 3.2 (Optimal excess risk in non-i.i.d. private FL).
Theorem 3.1 gives the first optimal rate for non-smooth heterogeneous FL under ISRL-DP. Note that the current best result for heterogeneous FL in Lowy and Razaviyayn (2023a) is suboptimal and holds only for the smooth case. One can combine the same smoothing technique above with the algorithm in Lowy and Razaviyayn (2023a) to obtain a risk bound of $O(1/\sqrt{nN}+(\sqrt{d\ln(1/\delta)}/{\varepsilon n\sqrt{N}})^{2/3})$ for the non-smooth case, which is again suboptimal.
Remark 3.3 (Improved communication complexity).
The communication complexity bound in Theorem 3.1 improves over the previous state-of-the-art result for an algorithm achieving optimal excess risk Lowy and Razaviyayn (2023a); recall (3). Moreover, the result of Lowy and Razaviyayn (2023a) assumed i.i.d. silo data, whereas our result holds for the non-i.i.d. case.
We have omitted the gradient complexity for this smoothing approach. This is mainly because the gradient computation of $f_{\beta}$ needs additional computation in the form of the prox operator. One may consider using an approximate prox operator to get a handle on the total gradient complexity as in Bassily et al. (2019). However, this approach would still introduce an additional $\Theta(n^{3})$ gradient complexity per step. Thus, a natural question is whether we can design a more computation-efficient algorithm for the non-smooth case — providing a motivation for our next algorithm.
3.2 Error-Optimal and Computationally Efficient ISRL-DP FL for Non-Smooth Losses
We propose another variation of Algorithm 1 that uses subgradients to handle the non-smooth case in a computationally efficient way: see Algorithm 4. Algorithm 4 follows the same structure as Algorithm 1: we iteratively solve a carefully chosen sequence of regularized ERM problems with a ISRL-DP solver and use localization. Compared to Algorithm 1 for the smooth case, Algorithm 4 does not use an accelerated solver (due to non-smoothness). Instead, we use ISRL-DP Minibatch Subgradient method (Algorithm 3) to solve the non-smooth strongly convex ERM problem in each phase of Algorithm 4. There are two key differences between our subroutine Algorithm 3 and the ISRL-DP MB-SGD of Lowy and Razaviyayn (2023a): (i) Instead of the gradient, a subgradient of the non-smooth objective is used in Line 6; (ii) A different and simpler averaging step in Line 11 is used for strongly convex non-smooth losses.
Algorithm 3 Noisy ISRL-DP MB-Subgradient Method
1:Datasets $X_{l}\in\mathcal{X}^{n}$ for $l\in[N]$, loss function $\hat{F}(w)=\frac{1}{nN}\sum_{l=1}^{N}\sum_{x\in X_{l}}f(w,x)$, constraint set $\mathcal{W}$, initial point $w_{0}$, privacy parameters $(\varepsilon,\delta)$, iteration count $R\in\mathbb{N}$, batch size $K\in[n]$, step sizes $\{\gamma_{r}\}_{r=0}^{R-1}$, initial point $w_{0}\in\mathcal{W}$.
2:for $r\in\{0,1,\ldots,R-1\}$ do
3:     for $l\in S_{r}$ in parallel do
4:         Server sends global model $w_{r}$ to silo $l$.
5:         Silo $l$ draws $K$ samples $x_{l,j}^{r}$ uniformly from $X_{i}$ (for $j\in[K]$) and noise $u_{i}\sim\mathcal{N}(0,\sigma^{2}\mathbf{I}_{d})$ for proper $\sigma^{2}$.
6:         Silo $l$ computes $\widetilde{g}_{r}^{l}:=\frac{1}{K}\sum_{j=1}^{K}g_{r,j}^{l}+u_{i}$ and sends to server, where $g_{r,j}^{l}\in\partial f(w_{r},x_{l,j}^{r})$ (subgradient).
7:     end for
8:     Server aggregates $\widetilde{g}_{r}:=\frac{1}{M_{r}}\sum_{l\in S_{r}}\widetilde{g}_{r}^{l}$.
9:     Server updates $w_{r+1}:=\Pi_{\mathcal{W}}[w_{r}-\gamma_{r}\widetilde{g}_{r}]$.
10:end for
11:Output: $\bar{w}_{R}=\frac{2}{R(R+1)}\sum_{r=1}^{R}rw_{r}$.
Algorithm 4 Localized ISRL-DP MB-Subgradient Method
1:Dataset $X_{l}\in\mathcal{X}^{n},\,l\in[N]$, constraint set $\mathcal{W}$, $\eta>0$, subroutine parameters (specified in Appendix) including batch size $K_{i}$, number of rounds $R_{i}$, noise parameters $\sigma_{i}$.
2:Choose any $w_{0}\in\mathcal{W}$.
3:Set $\tau=\lfloor\log_{2}n\rfloor$, $p=\max(\tfrac{1}{2}\log_{n}(M)+1,3)$.
4:for $i=1$ to $\tau$ do
5:     Set $\eta_{i}\!=\!\eta/2^{ip}$, $n_{i}\!=\!n/2^{i}$, $\lambda_{i}\!=\!1/(\eta_{i}n_{i})$, $D_{i}\!=\!2L/\lambda_{i}$.
6:     Each silo $l\in[N]$ draws disjoint batch $B_{i,l}$ of $n_{i}$ samples from $X_{l}$.
7:     Let $\hat{F}_{i}(w)=\frac{1}{n_{i}N}\sum_{l=1}^{N}\sum_{x_{l,j}\in B_{i,l}}f(w;x_{l,j})+\frac{\lambda_{i}}{2}\lVert w-w_{i-1}\rVert^{2}$.
8:     Call the $(\varepsilon,\delta)$-ISRL-DP Algorithm 3 with loss function $\hat{F}_{i}(w)$, data $X_{l}=B_{i,l}$, $R=R_{i}$, $K=K_{i}$, step sizes $\gamma_{r}=\frac{2}{\lambda_{i}(r+1)}$ for $r=0,1,\ldots,R_{i}-1$, initialization $w_{i-1}$, and constraint set $\mathcal{W}_{i}=\{w\in\mathcal{W}:\lVert w-w_{i-1}\rVert\leq D_{i}\}$. Let $w_{i}$ denote the output.
9:end for
10:return the last iterate $w_{\tau}$.
Theorem 3.4 (Non-smooth FL via subgradient).
Let $M=N$. Then, there exist parameter choices such that Algorithm 4 is $(\varepsilon,\delta)$-ISRL-DP and achieves the optimal excess population risk in (9). The communication complexity is
$\widetilde{O}\left(\min\left(nN,\frac{N\varepsilon^{2}n^{2}}{d}\right)+1\right).$
Assuming $\varepsilon=\Theta(1)$, the subgradient complexity is
$\widetilde{O}\Bigg{(}Nn+N^{2}\min\left(n,\frac{n^{2}}{d}\right)+N^{3/2}\min\left(n^{3/2},\frac{n^{2}}{\sqrt{d}}\right)\Bigg{)}.$
Remark 3.5 (Improved gradient complexity).
The above subgradient complexity improves over the state-of-the-art gradient complexity Lowy and Razaviyayn (2023a) for an ISRL-DP FL algorithm with optimal excess risk. Lowy and Razaviyayn (2023a) apply Nesterov smoothing to ISRL-DP MB-SGD. As discussed earlier, implementing the smoothing approach is computationally costly. Moreover, the results in Lowy and Razaviyayn (2023a) assume i.i.d. silo data. The subgradient complexity result in Theorem 3.4 also improves over the gradient complexity of Algorithm 1 with smoothing.
An alternative approach is to use the convolutional smoothing technique Kulkarni et al. (2021) to optimize non-smooth functions. For certain regimes, this approach improves the gradient complexity over the subgradient method, with a worse communication complexity compared to the Nesterov smoothing approach. See Appendix 9 for details.
Depending on the FL application, communication or computation may be more of a bottleneck. If communication efficiency is more of a priority, the smoothed version of Algorithm 1 should be used. On the other hand, if computational efficiency is more pressing, Algorithm 4 or Algorithm 6 is recommended.
4 Numerical Experiments
We validate our theoretical findings with numerical experiments on MNIST data. As shown in Figures 3 and 3, our algorithm consistently outperforms Lowy et al. (2023a). We use a similar experimental setup to Lowy et al. (2023a), as outlined below.
Task/model/data set.
We run binary logistic regression with heterogeneous MNIST data: each of the $N=25$ silos contains data corresponding to one odd digit class and one even digit class (e.g. (1, 0), (3, 2), etc.) and the goal is to classify each digit as odd or even. We randomly sample roughly $1/5$ of MNIST data to expedite the experiments. We borrow the code from Woodworth et al. (2020) to transform and preprocess MNIST data.
Preprocessing.
We preprocess MNIST data, flatten the images, and reduce the dimension to $d=50$ using PCA. We use an 80/20 train/test split, yielding a total of $n=1734$ training samples.
Our algorithm.
Localized ISRL-DP MB-SGD, which is a practical (non-accelerated) variant of our Algorithm 1: For simplicity and to expedite parameter search, we use vanilla MB-SGD in place of accelerated MB-SGD as our regularized ERM subsolver in our implementation of Algorithm 1.
Baseline.
We compare our algorithm against the One-pass ISRL-DP MB-SGD of Lowy et al. (2023a). Recall that Lowy et al. (2023a) did not provide theoretical guarantees for their multi-pass ISRL-DP MB-SGD with heterogeneous silos.
Hyperparameter tuning and evaluation.
We evaluate the algorithms across a range of privacy parameters $\varepsilon$ and fix $\delta=1/n^{2}$. For each algorithm and each setting of $\varepsilon$, we search a range of step sizes $\eta$.
Simulating unreliable communication.
In addition to the reliable communication setting where all silos communicate in each round $M=N=25$, we simulate unreliable communication by randomly selecting a subset of silos to communicate in each round. In each communication round, $M=18$ of the $N=25$ silos are chosen uniformly at random to communicate with the server.
Evaluation.
Each evaluation consists of 5 trials, each with different data due to random sampling. In each trial, for each parameter setting, we repeat 3 runs and choose the hyperparameters with the lowest average loss, and record the average test error as the test error. We plot the average test error and the standard deviation across trials.
Figure 2: Reliable Communication
Figure 3: Unreliable Communication
As shown in the plots, in both reliable and unreliable communication settings, our localized ISRL-DP MB-SGD algorithm outperforms the baseline one-pass ISRL-DP MB-SGD algorithm across all privacy parameters. Despite being an algorithm designed to achieve theoretical guarantees, our algorithm evidently performs well in practice.
5 Concluding Remarks and Open Questions
We have studied private federated learning in the absence of a trusted server. We characterized the minimax optimal excess risk bounds for heterogeneous ISRL-DP FL, answering an open question posed by Lowy and Razaviyayn (2023a). Further, our algorithms advanced the state-of-the-art in terms of communication and computational efficiency. For smooth losses, the communication complexity of our optimal algorithm matches the non-private lower bound.
To conclude, we discuss some open problems that arise from our work. (1) A rigorous proof of a ISRL-DP communication complexity lower bound. (2) Is there an optimal ISRL-DP algorithm with $O(nN)$ gradient complexity? A promising approach may be to combine Algorithm 1 with ISRL-DP variance-reduction. (Note that the gradient-efficient variance-reduced central DP algorithm of Zhang et al. (2022) uses output perturbation, which requires a trusted server.) (3) Is it possible to achieve both optimal communication complexity and optimal gradient complexity simultaneously with a single algorithm? Even in the simpler centralized setting ($N=1$), this question is open. (4) What are the optimal rates for ISRL-DP FL problems beyond convex and uniformly Lipschitz loss functions?
Acknowledgements
XZ is supported in part by NSF CNS-2153220 and CNS-2312835. CG, AL, and SW supported in part by NSF CCF-2023239 and CCF-2224213 and AFOSR FA9550-21-1-0084.
Impact Statement
In this research, we advance the field of federated learning (FL) by introducing algorithms that enhance privacy and communication efficiency in settings where trust in a central server is not assumed. Our work has significant implications for industries handling sensitive data, such as healthcare and finance, by offering an improved method for leveraging collective data while safeguarding individual privacy. However, the increased algorithmic complexity of these algorithms could limit their accessibility, especially for organizations with limited resources. Additionally, while our approach reduces the risk of data leakage, it does not entirely eliminate the possibility of misuse or unintended consequences, such as reinforcing existing biases in data. Future research should focus on refining these algorithms to make them more accessible and addressing potential ethical implications. By doing so, we aim to contribute positively to the development of safe and equitable data handling practices in various sectors of society and the economy.